Circular No. 8728

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2575 OPHIUCHI R. J. Rudy, D. K. Lynch, C. C. Venturini, S. Mazuk, D. Broadwater,

and M. Cowan, The Aerospace Corporation; R. C. Puetter, University of California at San Diego; and R. B. Perry, Langley Research Center, NASA, report 0.47- to 2.5-μm spectroscopy of V2575 Oph (cf. IAUC 8671, 8710) on June 14.34 UT using the VNIRIS spectrograph on the Lick 3-m telescope. The nova shows strong H I and very strong He I emission with He II weakly present. The object is heavily reddened, with E(B − V) = 1.5 as deduced from the O I lines. [O I] and [N II] emission lines are present at the optical wavelengths. The permitted emission lines display broad, doubled profiles with FHWM ∼ 1700 km/s. The profiles of the forbidden lines are narrower (1100 km/s) and are not doubled. There is no evidence for emission from dust. The following magnitudes for V2575 Oph were determined from the spectroscopy: V = 15.3, J = 10.7, H = 11.0, K = 10.1.

SUPERNOVAE 2006dm, 2006dn, 2006do Three additional supernovae have been discovered on unfiltered KAIT

images taken in the course of the Lick Observatory Supernova Search (cf. IAUC 8727), as reported by J. Schwehr and W. Li:

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006dm July 3.44 23^{h}41^{m}47.^{s}84 − 3^{o}40′08′′.4 17.3 7′′.6 E, 6′′.3 S
2006dn July 5.45 22 47 37.84 +39 52 46.8 17.8 4′′.5 E, 2′′.9 N
2006do July 5.48 23 32 03.08 −27 43 31.8 16.7 5′′.0 W, 2′′.4 N

Additional KAIT magnitudes for 2006dm in MCG –01-60-21: June 21.49 UT, [18.5; July 4.49, 16.9. SN 2006dm is a type-Ia supernova that was discovered a few days before maximum light (cf. CBET 569, 570). Additional KAIT magnitudes for 2006dn in UGC 12188: 2005 Dec. 17.10, [19.0; 2006 June 9.49, [18.5; 26.49, 18.7. SN 2006dn is a type-Ib supernova, discovered around a week past maximum (cf. CBET 573). Additional KAIT magnitudes for 2006do in ESO 470-G18: 2005 Dec. 14.09, [19.0; 2006 June 24.49, [18.0; July 6.48, 16.7. SN 2006do is a type-Ia supernova, discovered a few days before maximum (cf. CBET 573).

2006 July 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green